Graph each function
f(x)=|x + 2|; D = {-2,-1,0,1,2}

Answer by Edwin McCravy(20055)

f(x)=|x + 2|                  (x,y)
------------------------------------
f(-2)=|-2 + 2| = |0| = 0      (-2,0)
f(-1)=|-1 + 2| = |1| = 1      (-1,1)
f(0)=|0 + 2| = |2| = 2        (0,2)
f(1)=|1 + 2| = |3| = 3        (1,3) 
f(2)=|2 + 2| = |4| = 4        (2,4) 

plot them:



Edwin